OpenStudy (anonymous):

Determine a and b so that the vectors < 1,3a+b,2 > and < 2,−1,a+ b > are parallel.

OpenStudy (anonymous):

would i set up 3a+b=-1

OpenStudy (anonymous):

if they're parallel one should be a scalar multiple of the other. as such the second should be double the first. set up the equations in that way

OpenStudy (anonymous):

3a+b=-(1/2) a+b = 4

OpenStudy (anonymous):

where did you get the <-1/2> from

OpenStudy (anonymous):

the econd vector should be double the first

OpenStudy (anonymous):

i was thinking more along the lines of <1,3a+b,2>=S<2,-1,a+b>

OpenStudy (anonymous):

let u = <1, 3a+b, 2> and 2u = <2, 6a+2b, 4>, right? v= <2, -1, a+b> and if v = 2u then 6a+2b=-1 => 3a+b = -(1/2) a+b = 4

OpenStudy (anonymous):

ok i understand that but arent they looking for a +b individually or it doesnt matter

OpenStudy (anonymous):

yes and you have a system of 2 equations in 2 unknowns... solve using whatever method you prefer.

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

you're welcome... does it all make sense now?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

cool!